Label Corp. recorded sales of $2,235,106 for the most recent year. The company's breakeven sales point (in dollars) is $1,750,000, and its margin of safety ratio (MOS%) at the current sales level is 26%. What sales (in dollars) would be needed to increase the company's MOS% to 32%?

Multiple Choice

  • $2,189,239.

  • $2,573,529.

  • $2,844,879.

  • $2,936,289.
Answer #1

Margin of safety = (Sales-Break even sales) /Sales

Let sales be x

(x - 1,750,000)/x = 0.32

x - 1,750,000 = 0.32x

x - 0.32x = 1750,000

x = 1,750,000/0.68

= 2,573,529

Option B